Find the range of 1.f(x,y,z)= z/(x^2-y^2)
2. f(x,y)= x sin(1/x)+ y sin(1/y)
3. f(x,y,z)= 1/(√ (4-x^2-y^2-z^2)
1
Expert's answer
2021-07-27T09:17:10-0400

Answer:-


1.


"x^2-y^2\not=0"

If "z\geq 0, x^2>y^2," then "f(x, y,z)\geq0."

If "z\geq 0, x^2<y^2," then "f(x, y,z)\leq0."

If "z\leq 0, x^2>y^2," then "f(x, y,z)\leq0."

If "z\leq 0, x^2<y^2," then "f(x, y,z)\geq0."

Range: "(-\infin, \infin)"


2. Let "u(x)=x\sin(\dfrac{1}{x})"

If "x\to\pm \infin," then "u(x)\to 1^{-}"



"u'=\sin(\dfrac{1}{x})-\dfrac{1}{x}\cos(\dfrac{1}{x})""u'=0=>\sin(\dfrac{1}{x})-\dfrac{1}{x}\cos(\dfrac{1}{x})=0""x_1=-0.22255, x_2=0.22255""u(-0.22255)=u(0.22255)\approx-0.21723""-0.21723\leq u<1, x\not=0""-0.21723-0.21723\approx-0.4345"



Range: "[-0.4345, 2)"


3.


"4-x^2-y^2-z^2>0""0\leq x^2+y^2+z^2<4"

Then



"0<\sqrt{4-x^2-y^2-z^2}\leq2"

Range: "\bigg[\dfrac{1}{2}, \infin\bigg)"